### v2

# Neural Signals Generate Clinical Notes in the Wild

Abstract
Generating clinical reports that summarize abnormal patterns, diagnostic findings, and clinical interpretations from long-term EEG recordings remains labor-intensive. We curate a large-scale clinical EEG dataset with 9,9229{,}922 reports paired with approximately 11,00011{,}000 hours of EEG recordings from 9,0489{,}048 patients. We therefore develop CELM, the first clinical EEG-to-Language foundation model capable of summarizing long-duration, variable-length EEG recordings and performing end-to-end clinical report generation at multiple scales, including recording description, background activity, epileptiform abnormalities, events/seizures, and impressions. Experimental results show that, with patient history supervision, our method achieves $70\%$$95\%$ average relative improvements in standard generation metrics (e.g., ROUGE-1 and METEOR) from 0.20.20.30.3 to 0.40.40.60.6. In the zero-shot setting without patient history, CELM attains generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. CELM integrates pretrained EEG foundation models with language models to enable scalable multimodal learning. We release our model and benchmark construction pipeline at https://github.com/Jathurshan0330/CELM.
Machine Learning, ICML
1 Introduction
Electroencephalography (EEG) records long-term, real-time neuronal activity with millisecond-level temporal resolution and is widely used in hospitals for neurological diagnosis, such as epilepsy (Kotoge et al., 2025) and sleep health (Yang et al., 2023). However, writing clinical reports to summarize diagnostic findings over long-term EEG recordings remains a labor-intensive process. Typically, neurologists visually inspect EEG signals to assess brain activity and then compose clinical reports that document abnormal patterns (i.e., EEG phenotypes) and their clinical interpretations (Tatum IV et al., 2016). Moreover, this task requires substantial domain knowledge and extensive clinical experience, which in turn incur long-term training costs and require continual updates as new EEG phenotypes or waveforms are identified (Biswal et al., 2019).
To support this profound clinical practice, several automated clinical EEG report generation models based on advanced deep learning techniques have been proposed (Biswal et al., 2019; Lee et al., 2023; Yin and Shin, 2025). These methods aim to encode EEG recordings into latent representations that capture temporal structure and clinically relevant phenotypes, which are then decoded into clinical reports. However, existing methods remain limited in several aspects. ① Conceptually, some works formulate report generation as a phenotype classification or retrieval problem followed by text decoding. They do not model EEG-to-report generation as an end-to-end learning task, and the misalignment between classification and generation objectives does not guarantee optimal report generation. ② Methodologically, these methods operate on short EEG segments and fixed-context templates. They fail to model long-term temporal context and global diagnostic reasoning required for interpreting long-duration EEG data. Moreover, fixed-context templates prevent multi-granularity report generation beyond predefined categories. ③ Practically, these methods are developed as task-specific models, each tailored to a narrow reporting objective (e.g., impression-only reports) (Lee et al., 2023). In contrast, clinical practice requires multi-level reporting, where neurologists routinely construct overall EEG summaries and section-wise outputs, including impressions, interpretations, and event or seizure annotations.
Present work. We introduce CELM , which, to the best of our knowledge, is the first clinical EEG-to-language foundation model (ELM) capable of generating EEG reports at multiple scales. We train and evaluate CELM on 9,9229{,}922 clinical reports paired with approximately 11,00011{,}000 hours of EEG recordings from 9,0489{,}048 patients, spanning both clinical context–conditioned and zero-context settings, and covering key report sections including EEG description, background activity, epileptiform abnormalities, events/seizures, and impressions/interpretations.
Our motivation stems from recent advances in multimodal foundation models that enable text generation from non-text modalities (e.g., images and speech) (Bordes et al., 2024; Shen et al., 2025). Recent studies have begun to incorporate EEG signals into language-based modeling frameworks (Ndir et al., 2025; Gijsen and Ritter, 2025). However, they primarily leverage text supervision to enhance EEG classification and do not support EEG-to-text report generation.
To tackle aforementioned limitations, the contributions of this paper lie in: ❶ Conceptually, we formulate neurological clinical report generation as an EEG-to-Language foundation modeling problem. CELM enables end-to-end training and generation by directly translating raw EEG recordings into clinical text, without relying on intermediate phenotype classification or template-based pipelines. ❷ Methodologically, we introduce the first EEG-to-language multimodal framework, comprising three core components: (i) epoch-aggregated tokenization that produces compact representations from variable-length recordings, (ii) sequence-aware alignment that captures inter-epoch temporal dependencies, and (iii) prompt fusion for conditional report generation. Importantly, our model design harnesses the representation capabilities of pretrained EEG foundation models and LLMs while enabling flexible generation across multiple clinical settings. ❸ Practically, we introduce the first ELM benchmark construction pipeline based on the Harvard Electroencephalography Database (Sun et al., 2025; Zafar et al., 2025), which contains approximately $12\mathrm{k}$ clinical reports paired with about $100\mathrm{k}$ hours of EEG recordings from $10k$ patients. Rather than requiring access to the raw, unstructured 200200TB corpus and associated metadata, our pipeline enables efficient filtering of EEG recordings matched to reports. We release all source code to support further evaluation and development.
Experimental results show that CELM consistently outperforms all baselines across all settings. With patient history, it achieves $70\%$$95\%$ average relative improvements, with standard generation metrics (e.g., METEOR (Banerjee and Lavie, 2005) and ROUGE-1 (Lin, 2004)) from 0.20.20.30.3 to 0.40.40.60.6 across diverse cases. In the zero-context setting without patient history, CELM achieves generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. We provide multi-perspective case analyses to shed light on future ELM development.
2 Related Work
EEG-to-Language Modeling.
Existing work at the intersection of EEG and natural language can be broadly categorized into two paradigms: (1) EEG-to-language decoding and (2) text-enhanced EEG representation learning. The goal of EEG-to-language decoding is to reconstruct textual content from concurrent EEG recordings of subjects reading or imagining speech (Herff et al., 2015; Wang and Ji, 2022). This line of work spans both invasive approaches using electrocorticography (ECoG) (Makin et al., 2020; Willett et al., 2023) and non-invasive methods based on scalp EEG (Wang and Ji, 2022; Duan et al., 2023; Zhou et al., 2024). However, these methods assume precise EEG–text alignment, whereas clinical EEG consists of heterogeneous events embedded in long, continuous recordings spanning hours to days (Sun et al., 2025; Zafar et al., 2025). The second paradigm leverages clinical notes as auxiliary supervision to enhance EEG representation learning, rather than decoding language directly from neural activity. Inspired by vision–language pretraining frameworks (Radford et al., 2021), a recent work  (Ndir et al., 2025) aligns EEG data with clinical report text in a shared feature space. However, it remains focused on discriminative objectives and does not support automated clinical report generation.
Clinical EEG Report Generation. An early attempt at EEG report generation, EEGtoText (Biswal et al., 2019), proposed a two-stage pipeline that first classifies EEG phenotypes, and then generates report text via a text decoder conditioned on the classified labels. However, this relies on intermediate phenotyping as a bottleneck, limiting the capabilities to capture nuanced clinical findings beyond predefined categories. While some attempts jointly learn EEG encoders and text decoders (Lee et al., 2023), they rely on fixed segmentation and template-based generation (Yin and Shin, 2025). As a result, these methods fall short of end-to-end clinical report generation from long-duration EEG recordings. We address the above challenges by introducing the first family of EEG–Language foundation models.
EEG Foundation Models. Recent years have witnessed rapid advances in EEG foundation models. BENDR (Kostas et al., 2021), BIOT (Yang et al., 2023), LaBraM (Jiang et al., 2024), TFM-Tokenizer (Pradeepkumar et al., 2025), EEGPT (Wang et al., 2024a), LUNA (Döner et al., 2025), REVE (Ouahidi et al., 2025), and CBraMod (Wang et al., 2024b) introduce increasingly scalable tokenization and representation learning frameworks that achieve strong transfer on diverse EEG tasks. These models are predominantly encoder-only and are optimized for classification tasks, leaving generative objectives largely unexplored. Our CELM is designed for clinical EEG report generation, and notably, it is fully compatible with existing foundation models, which can be directly leveraged as EEG encoders.
3 EEG-Report Benchmark Construction
Figure 1: Overview of our framework. (a) EEG–Report benchmark construction pipeline, including clinical report structuring (Section 3.1), matching reports to EEG sessions, and examples of standardized report sections. (b) The proposed Clinical EEG Language Model (CELM) comprises (b.1) Epoch-Aggregated Tokenization, (b.2) Sequence-Aware Alignment, and (b.3) Prompt Fusion and Generation.
Currently, no publicly available benchmark provides structured EEG–report pairs for training and evaluating EEG-to-Language Models (ELMs). In this work, we leverage two resources from the Brain Data Science Platform: (1) the Harvard Electroencephalography Database v4.1 (Zafar et al., 2025; Sun et al., 2025), which contains approximately $10k$ clinical reports paired with about $100k$ hours of EEG recordings from 10,88610{,}886 patients across multiple hospital sites, and (2) the Electronic Health Records (EHR) Repository with corresponding unstructured neurology reports. The EHR repository links reports to one or more EEG sessions via temporal alignment between EDF start times and report timestamps. Therefore, we construct an EEG–report benchmark using a dedicated processing pipeline (Figure 1a). Specifically, Section 3.1 describes how neurology reports are extracted and organized based on whether their content is inferable from EEG data. Section 3.2 details the preprocessing and standardization of the matched EEG recordings.
3.1 Clinical Report Structuring
Neurology reports can be broadly decomposed into two types of sections: (1) Clinical Context, which contains information such as patient history and monitoring indications that cannot be inferred from EEG signals, and (2) EEG Findings, which describe EEG observations and clinical interpretations that are directly inferable from EEG recordings. This distinction is critical for evaluating clinical ELMs, as sections within EEG Findings provide the appropriate ground truth for benchmarking a model’s ability to generate clinical narratives from EEG signals.
A key challenge lies in reliably extracting and structuring these sections without information loss. Substantial variability in reporting styles and formatting makes manual extraction infeasible at scale. We therefore propose a three-stage automated pipeline to extract and structure reports.
Section Detection: first, we curate a comprehensive list of candidate section headers under each category and detect their presence using string matching.
Section-wise Extraction: second, for each detected section, we employ a local LLM (Meta-Llama-3-8B-Instruct) to extract the corresponding text using simple copy-based prompts. This detect-then-segment strategy avoids hallucinations and cross-section contamination that arise when extracting multiple sections jointly. Local deployment is required, as data use agreements prohibit sending clinical reports to proprietary cloud-based LLMs.
Canonical Normalization: finally, we standardize the extracted sections into canonical categories, including EEG description/details, impressions/interpretations, background activity, epileptiform abnormalities, and events/seizures.
3.2 EEG Preprocessing
Matched EEG recordings are preprocessed using a standard pipeline. Signals are first band-pass filtered between 0.1–75 Hz and notch filtered at 60 Hz to remove power-line interference. The recordings are then resampled to 200 Hz to ensure compatibility with existing EEG-only foundation models. Each recordings are segmented into non-overlapping 10-second windows, which is the typical temporal context used by current EEG foundation architectures. Finally, EEG channels are standardized to a 22-channel montage following the International 10-20 system. While the dataset contains reports matched to multiple EEG sessions, we filter to samples with single-session matches for training, validation, and evaluation. We then perform patient-level splitting into 60/20/20 training, validation, and test sets to ensure no patient overlap across splits, thereby preventing data leakage and producing a structured EEG–report benchmark.
4 CELM: Clinical EEG Language Model
This section introduces our CELM , consisting of three key components: Epoch-Aggregated Tokenization, Sequence-Aware Alignment, and Prompt Fusion and Generation. Unlike classification-oriented EEG models that learn discriminative representations for categorical labels, clinical ELMs must identify and aggregate clinically relevant information from EEG recordings to support report generation at multiple levels. To this end, Section 4.1 discusses the unique challenges of clinical ELMs and provides an overview of the model architecture and forward process. Sections 4.2, 4.3, and 4.4 then present the details of the three components.
4.1 Architecture and Forward Process
Clinical EEG introduces three unique challenges that prevent the direct adoption of existing multimodal paradigms, such as dictionary-based alignment or contrastive learning with fixed-length epochs (Ndir et al., 2025). Here, we design CELM with three corresponding steps:
Epoch-Aggregated Tokenization addresses the extreme temporal scale of clinical EEG. EEG sessions span hours of high-frequency, multi-channel signals, resulting in at least $\sim 31.7\mathrm{M}+$ data points per recording, far exceeding the context limits of modern LLMs (Grattafiori et al., 2024; Team et al., 2025; Yang et al., 2025).
Sequence-Aware Alignment mitigates the loss of temporal context during alignment. EEG signals are inherently sequential, and naive projection into the LLM embedding space fails to preserve long-range temporal dependencies required for clinical interpretation.
Prompt Fusion and Generation addresses weak EEG–text correspondence. Clinical reports aggregate findings over entire recordings without explicit temporal grounding, requiring the model to synthesize coherent clinical narratives from distributed EEG evidence.
Overall, given an input EEG recording, CELM processes the signal through above three stages. The recording is first transformed into compact epoch-level tokens via Epoch-Aggregated Tokenization, which reduces the extreme temporal scale. These tokens are then projected into the language model space through Sequence-Aware Alignment, enabling long-range temporal dependencies to be maintained across epochs. Finally, Prompt Fusion and Generation conditions the language model on the aligned EEG representations and optional clinical context to generate structured reports.
4.2 Epoch-Aggregated Tokenization
The fundamental bottleneck in developing ELMs is representing long EEG recordings within the LLM context constraints. A typical two-hour clinical EEG recording at $200$ Hz across $22$ channels yields $7,200\times 200\times 22\approx 31.7$M data points, far exceeding current LLM capacities, making raw sample-level tokenization infeasible. A common alternative is mini-window tokenization, leveraging pretrained EEG encoders that typically tokenize 1-second segments. However, naively applying this granularity to full recordings produces $\approx 158$K tokens per session, still well beyond the context limits of most local LLMs (Grattafiori et al., 2024; Team et al., 2025; Yang et al., 2025). To overcome this limitation while harnessing pretrained EEG representations, we propose Epoch-Aggregated Tokenization (see Figure 1b.1).
Given an EEG recording session $\mathbf{X}\in\mathbb{R}^{N\times C\times T}$, where $N,C$, and $T$ denote the number of epochs, channels, and time points per epoch (typically 10 seconds) respectively, we first tokenize mini $1$s windows within each epoch using a pretrained EEG encoder. The resulting representations are then aggregated across mini-windows and channels to produce a single token per epoch, yielding a compact sequence of epoch tokens $\mathbf{E}_{\text{eeg}}\in\mathbb{R}^{N\times D_{\text{eeg}}}$, where $D_{\text{eeg}}$ is the embedding dimension. The aggregation strategy is encoder-dependent. For example, CBraMod applies pooling (Wang et al., 2024b), whereas LaBraM (Jiang et al., 2024) and TFM-Tokenizer (Pradeepkumar et al., 2025) uses [CLS] pooling. This tokenization compresses the tokens by up to $C\times T$ ($\sim 220\times$ compared to mini-window tokenization), enabling EEG recordings to fit within LLM context limits.
4.3 Sequence-Aware Alignment
To enable EEG-conditioned clinical report generation and leverage LLMs’ generative capabilities, EEG tokens must be aligned with the text embedding space of LLMs. In the vision-language setting, models such as LLaVA achieve this via a simple linear projection (Liu et al., 2023). However, this paradigm does not generalize well to EEG (Section 5.4) due to its inherent sequential nature, unlike static images. EEG signals exhibit long-range temporal dynamics across epochs that must be preserved during alignment. As a result, we face two key challenges: modeling long temporal dependencies and scalability for very long sequences. These challenges motivate the design of two alignment strategies in the proposed Sequence-Aware Alignment module (see Figure 1b.2): (1) Sequence Context Alignment, which preserves the full temporal sequence, and (2) Sequence Context Compression, which compresses the sequence into a fixed-length latent representation for memory-efficient alignment.
Sequence Context Alignment (SCA).
Given a sequence of epoch tokens $\mathbf{E}_{\text{eeg}}$, we augment the sequence with learnable position encodings and [SESSION_SEP] tokens to differentiate EEG sessions. A lightweight linear-attention transformer (Wang et al., 2020; Katharopoulos et al., 2020) is then applied to capture the temporal structure across epochs. The resulting representations are subsequently projected through a linear layer into the language embedding space, yielding $\mathbf{H}_{\text{eeg}}\in\mathbb{R}^{N\times D_{\text{llm}}}$, where $D_{\text{llm}}$ denotes the LLM embedding dimension.
Sequence Context Compression (SCC).
In an attempt to tackle the memory constraints and improve scalability for long EEG sequences, we explore a compression strategy inspired by Perceiver architectures (Jaegle et al., 2021b, a). After encoding temporal structure among epoch tokens, a fixed set of learnable query tokens $\mathbf{Q}\in\mathbb{R}^{L\times D_{\text{eeg}}}(L<N)$ attend to the full sequence of epoch tokens via cross-attention. This compresses the variable-length EEG sequence into a fixed number of tokens, which are subsequently projected into the language embedding space, enabling efficient alignment for recordings of arbitrary duration.
4.4 Prompt Fusion and Generation
CELM generates clinical reports by conditioning a local LLM on EEG-derived representations jointly with clinical context when available (see Figure 1b.3). Given projected EEG tokens $\mathbf{H}_{\text{eeg}}$, we augment the sequence with special tokens [EEG_START] and [EEG_END] to distinguish EEG representations from text inputs. A task-specific prompt that specifies the target report sections, together with optional clinical context (e.g., reason for study or clinical history), is tokenized using the LLM tokenizer to obtain text embeddings $\mathbf{H}_{\text{prompt}}$. The final input sequence to the LLM is formed by concatenation,
$\mathbf{H}_{\text{in}}=[\texttt{[EEG\_START]};\mathbf{H}_{\text{eeg}};\texttt{[EEG\_END]};\mathbf{H}_{\text{text}}]$ (1)
which is processed autoregressively to generate a structured EEG report text. We employ instruction tuned Qwen3-4B (Yang et al., 2025) as our base LLM, CBraMod (Wang et al., 2024b) as our EEG encoder and perform supervised fine-tuning using the next-token prediction objective:
$\mathcal{L}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid H_{\text{input}},y_{<t})$ (2)
During training, both the EEG encoder and LLM remain frozen, only the Sequence-Aware Alignment module is updated, enabling efficient adaptation while preserving pretrained representations.
5 Experiments and Results
Table 1: Report generation performance comparison on samples with clinical context across two hospital sites (S0001 and S0002). We evaluate BLEU-1, ROUGE-1, and METEOR under two input settings: (i) Unimodal + Text Only Input, where language models generate reports solely from clinical context text, and (ii) Unimodal + Text + EEG Features Input, where EEG-derived features are additionally provided. The table compares multiple strong baselines, including general-purpose and medical LLMs, against our CELM . The best results are highlighted in orange, and the best baseline per category is highlighted in blue. Relative improvements are also reported.
[…]
Datasets.
The EEG-Report benchmark is constructed using data from two hospital sites in the Harvard Electroencephalography Database v4.1 (Zafar et al., 2025; Sun et al., 2025): Massachusetts General Hospital (MGH, S0001) and Brigham and Women’s Hospital (BWH, S0002). In total, through our pipeline, we curated $12,290$ clinical EEG reports matched to one or more EEG sessions from $10,886$ patients. For model training and evaluation, we restrict our analysis to reports paired with a single EEG session.
S0001: This site includes diverse EEG visit types (routine EEG, epilepsy monitoring unit (EMU), portable EEG, complex EEG). We filtered only routine EEG and EMU samples paired with single sessions, excluding long recordings exceeding $10,000$ seconds to prevent memory issues during training. We curated 5,0495{,}049 reports from 4,6694{,}669 patients.
S0002: From the available visit types (EEG, long-term monitoring, portable EEG), we retained only standard EEG visits with single corresponding sessions, yielding $4,873$ reports from $4,379$ patients.
For both sites, we performed patient-wise 60/20/20 splits into training, validation, and test sets. Additional dataset details are provided in Appendix A.
Baselines.
As no existing EEG multimodal models with generative capabilities are available for EEG report generation, we establish baselines by augmenting local LLMs with EEG-derived inputs. We evaluate two settings: (1) Unimodal + text-only input, where the LLMs receive only clinical context as input without EEG data, providing a lower bound that quantifies hallucination and measures performance when the model does not attend to neural signals, and (2) Unimodal + text and EEG features input, where we augment the clinical context with channel-wise spectral band power features (delta, theta, alpha, beta, gamma) extracted from each EEG session, addressing context length limitation that prevents providing raw EEG arrays directly. We evaluate LLMs from three model families, including LLaMA-3 (Grattafiori et al., 2024), Gemma-3 (Team et al., 2025), and Qwen-3 (Yang et al., 2025), across multiple parameter scales. Additionally, we include MedGemma (Sellergren et al., 2025) as a medical LLM baseline. Additional implementation details are provided in Appendix B. We report smoothed BLEU-1 (Papineni et al., 2002), ROUGE-1 (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) scores to evaluate report generation performance. Comprehensive results and metrics are provided in Appendix D.
5.1 Report Generation Performance Comparison
Table 1 compares CELM against unimodal baselines using either text-only or text with handcrafted EEG features as input, evaluated on test samples containing patient history. We report results for two CELM variants: CELM-SCC, which uses sequence context compression to compress and produce a fixed number of projected EEG tokens, and CELM , which uses non-compressed EEG representations via sequence context alignment. When EEG features are provided, most baseline LLMs show modest improvements over text-only inputs on S0002 ($0.1912\rightarrow 0.2483$), except LLaMA-3.1-8B, which fails due to its 8K token context limit. However, on S0001, performance degrades across baselines due to greater report complexity introduced by multiple sections.
Both CELM variants substantially outperform all baselines across all metrics, with CELM achieving the best results. On average, CELM yields a two-fold improvement over the strongest baseline across both sites and all evaluation metrics. For example, on the ROUGE score for S0002, performance improves from $0.2886$ to $0.6408$ relative to the strongest baseline. Although CELM-SCC outperforms the baselines, its performance gap relative to the non-compressed model ($0.4487$ vs $0.6408$) underscores an important open challenge. This leads to an important future direction of developing efficient compression strategies that preserve clinically relevant information from long EEG sequences while remaining within LLM context constraints.
Table 2: Zero-context report generation performance on S0002. The best results are highlighted in orange, and the best baseline is highlighted in blue. Relative improvements are reported with respect to the strongest baseline under each setting.
[…]
Figure 2: Section-wise comparison of report generation performance between CELM and the best-performing baselines from different LLM families.
Figure 3: (a) Report generation performance of different projector variants. (b) Training dynamics of each variant, including training loss, validation loss, and perplexity curves.
Figure 4: Examples of generated reports on S0002. Comparisons between the unimodal baseline (text + EEG features), the linear-projector alignment variant, CELM, and the ground-truth reports. (a) EEG description/details; (b) Impression/interpretation.
5.2 Zero-Context Report Generation
To assess our model’s ability to generate clinical reports solely from EEG signals, without relying on clinical context, we define a zero-context report generation task in which only EEG data is provided as input. This experiment also serves as a test of potential language-model bias, in which reports could be generated by inferring plausible clinical context while ignoring EEG inputs. We evaluate this task on a subset of the S0002 dataset containing reports without patient clinical context, and the results are summarized in Table 2. CELM significantly outperforms all baselines across all metrics, demonstrating that it effectively extracts and utilizes clinically relevant information from EEG recordings.
5.3 Performance Analysis by Report Section
Clinical EEG reports comprise multiple sections, each serving a distinct purpose: EEG description/details provides a detailed narrative of observed waveforms and patterns, impression/interpretation summarizes the clinical significance of findings, background activity characterizes baseline rhythms, and events/seizures documents seizure episodes. To evaluate whether CELM can reliably generate these diverse sections, we conduct a section-wise analysis on S0001, which exhibits richer section diversity compared to S0002 (where reports predominantly contain only EEG description/details; see Figure 5 in Appendix). Figure 2 presents ROUGE-1 scores, comparing CELM against the best-performing unimodal baseline (text + EEG features) from each LLM family. CELM achieves the highest performance in 6 out of 7 report sections. Performance degrades in the interictal epileptiform abnormalities section, highlighting a key limitation and a challenge for ELMs in modeling rare and clinically complex events, an important direction for future work. Detailed results are provided in Table 12.
5.4 Alignment Module Ablation
The alignment module is a critical component of EEG-language models. In our initial experiments, we found that naive linear projection into the LLM embedding space is insufficient for the EEG domain. In this section, we present an ablation study comparing different alignment strategies. We evaluate two baseline alignment modules: (1) Linear projector, which directly maps EEG embeddings to the LLM embedding space using a single linear layer, and (2) Perceiver projector, which employs learnable query tokens that attend to EEG sequences via cross-attention followed by feedforward and projection. We compare these against our proposed SCA and SCC projectors, described in Section 4.3. Results are summarized in Figure 3.
As shown in Figure 3a, the SCA projector consistently achieves the best performance across all evaluation metrics. Analysis of training dynamics in Figure 3b further highlights the importance of modeling inter-epoch temporal dependencies prior to alignment. While the Linear and Perceiver projectors’ training losses decrease, their validation losses and perplexities increase, indicating overfitting. In contrast, both SCA and SCC exhibit more stable convergence, with lower validation loss and perplexity over time, and SCA converges faster across all variants. Notably, SCA achieves lower validation perplexity than other variants, indicating substantially higher confidence in token generation. Overall, this study confirms that modeling temporal dependencies between EEG epochs before projection is essential for effective EEG-to-language alignment.
5.5 Qualitative Analysis
Figure 4 presents qualitative comparisons on S0002 between the best-performing unimodal baseline (text with EEG features), the linear projector variant, and CELM , alongside ground truth reports. Figure 4a shows the EEG description/details section, while Figure 4b shows the impression/interpretation section. Across these examples, CELM produces reports that are more closely aligned with the ground truth, correctly identifying clinically relevant findings such as moderate bilateral beta activity, intermittent drowsiness, and bilateral slowing in the theta and delta ranges. In contrast, the linear projector and unimodal baselines fail to capture these. In the impression section, CELM identifies the recording as abnormal, whereas the linear projector variant incorrectly predicts it as normal. These results demonstrate the promise of ELMs in learning to generate clinical reports from unstructured notes paired with long EEG recordings. At the same time, they underscore the need for more rigorous and standardized benchmarking to assess these models. Additional examples spanning both sites, diverse report sections, and varying performance levels are provided in Appendix D.7 to further illustrate model strengths and failure modes.
6 Discussion and Implication
In this work, we demonstrate the feasibility of end-to-end clinical EEG–language modeling for generating structured reports directly from long-duration EEG recordings. The strong performance gains over competitive baselines highlight the potential of ELMs for clinical report generation. However, several limitations remain. ☞Evaluation limitations. ELM development is constrained by the lack of rigorous benchmarks and clinically grounded evaluation protocols, as common text generation metrics mainly capture lexical similarity rather than clinical correctness. ☞Memory scalability. Memory remains a key bottleneck. While our approach supports EEG recordings of up to approximately 3 hours, further advances in memory-efficient representations are needed to scale to longer recordings. ☞Human-in-the-loop potential. ELMs offer opportunities for human-in-the-loop workflows, where clinicians guide generation via targeted prompts instead of fixed templates, enabling better alignment with clinical intent.
7 Conclusion
We introduced CELM, the first clinical EEG-to-language foundation model for automated report generation from long-duration EEG recordings. CELM addresses three core challenges in this setting: (1) representing hour-scale EEG within LLM context limits via Epoch-Aggregated Tokenization, (2) preserving long-range temporal dependencies through Sequence-Aware Alignment, and (3) enabling flexible, multi-scale report generation via Prompt Fusion. We further presented a scalable EEG–Report benchmark construction pipeline and conducted extensive evaluations across hospital sites, report sections, and zero-context settings. Empirical results show that CELM consistently outperforms strong baselines. Overall, this work establishes EEG-to-language modeling as a distinct and promising research direction at the intersection of multimodal learning and clinical neurophysiology, and we hope that CELM and the accompanying benchmark construction pipeline will facilitate future advances in long-context EEG modeling and clinically grounded EEG language systems.
8 Impact Statement
This work advances machine learning research by introducing the first clinical EEG–language model that integrates long-duration clinical EEG signals with large language models to generate structured clinical reports. The proposed approach has the potential to support clinical workflows; however, this work is intended as a research contribution rather than a deployable clinical system. Any real-world application would require extensive validation, regulatory approval, and careful oversight by medical professionals. All data used in this study were deidentified by the dataset providers to protect participants’ protected health information (PHI). In addition, we strictly adhered to the data-use agreements and conducted all experiments using local large-language models.
Appendix
Appendix A Details of EEG-Report Benchmark
This section provides detailed information on the EEG–Report benchmark constructed in our study. Table 3 summarizes key statistics, including the number of EEG–report pairs, patient counts, and total EEG recording duration for each site. Figure 5 further illustrates dataset distributions across EEG session duration, patient demographics (age and gender), and EEG report sections.
Table 3: Statistics of the EEG–Report benchmark.
Site # of Paired EEG & Reports # of Patients Total Duration (hrs)
S0001 5049 4669 6639
S0002 4873 4379 4367
Figure 5: Dataset statistics of the filtered and constructed EEG-Report Benchmark used in our study
Appendix B Additional Experiment Details
B.1 Band Power Features
This section describes the extraction of handcrafted EEG features used in the unimodal text + EEG feature baselines. To construct EEG features that fit current context-length constraints, we extract band-power features from EEG recordings. For each 10-second multichannel EEG segment, we compute channel-wise band power using Welch’s method by first estimating the power spectral density and then integrating it within five EEG frequency bands: delta (0.5–4 Hz), theta (4–8 Hz), alpha (8–13 Hz), beta (13–30 Hz), and gamma (30–80 Hz). To further reduce input sequence length for LLMs with limited context windows, we optionally pool consecutive epochs by concatenating their time axes prior to spectral estimation, yielding a more compact representation. The resulting band power values are formatted as structured text and provided as input to the LLMs (see example in Figure 7).
B.2 Evaluation Metrics
We evaluate report generation performance using a comprehensive set of natural language generation metrics to evaluate different aspects of generation quality. In the main manuscript, we report smoothed BLEU-1, ROUGE-1, and METEOR scores. In Appendix D, we additionally report smoothed BLEU-4, ROUGE-2, ROUGE-L, and ROUGE-LSUM. We further include per-sample score distributions for all metrics in Appendix D.5.
B.3 Baselines and Prompts
In this section, we provide additional details on the baseline models and the prompts used in our experiments. Table 4 summarizes all LLMs considered in our study along with their respective context lengths [^1] Context length information is extracted from their respective technical reports and Hugging Face https://huggingface.co/, further underscoring the practical challenges of directly conditioning LLMs on long-duration EEG recordings. Figures 6 and 7 present the prompts used for the unimodal text-only and unimodal text + EEG feature baselines, respectively. For clarity, we provide examples for certain prompt sections using bounding boxes.
Table 4: Summary of LLM Baselines
LLM Family Context Length
Gemma-3-1b-it Gemma 3 32K
Gemma-3n-E2B-it Gemma 3 32K
Gemma-3n-E4B-it Gemma 3 32K
Gemma-3-4b-it Gemma 3 128K
Medgemma-4b-it Gemma 128K
Llama-3.2-1B-Instruct Llama-3.2 128K
Llama-3.2-3B-Instruct Llama-3.2 128K
Llama-3.1-8B-Instruct Llama-3.1 128K
Meta-Llama-3-8B-Instruct Llama-3 8K
Qwen3-1.7B Qwen 3 32K
Qwen3-4B-Instruct-2507 Qwen 3 256K
Unimodal + Text Only Prompt You are an expert clinical neurophysiologist specializing in EEG interpretation and clinical report generation. TASK
Your task is to generate the specified sections of a formal clinical EEG report using only the provided:
[…]
Return ONLY the following JSON structure, with no preamble, explanation, or markdown:
[…]
[SECTION_NAMES]
[…]
e.g. [’EEG DESCRIPTION/DETAILS’] PATIENT HISTORY AND EEG DESCRIPTION
[PATIENT_HISTORY_AND_EEG_DESCRIPTION]
[…]
e.g. age: 77.0, gender: Female, indication: ***** y.o. female with history of afib, l mca stroke, and headaches, presenting with episodes of altered consciousness concerning for seizure vs. syncope. none pertinent medications: keppra, neurontin, seroquel. none Now generate the EEG report.
Figure 6: Prompt for unimodal + text only baselines.
[…]
Your task is to generate the specified sections of a formal clinical EEG report using only the provided:
[…]
Return ONLY the following JSON structure, with no preamble, explanation, or markdown:
[…]
[SECTION_NAMES]
[…]
e.g. [’EEG DESCRIPTION/DETAILS’] PATIENT HISTORY AND EEG DESCRIPTION
[PATIENT_HISTORY_AND_EEG_DESCRIPTION]
[…]
e.g. age: 77.0, gender: Female, indication: ***** y.o. female with history of afib, l mca stroke, and headaches, presenting with episodes of altered consciousness concerning for seizure vs. syncope. none pertinent medications: keppra, neurontin, seroquel. none Now generate the EEG report.
Figure 7: Prompt for unimodal + text and EEG features as input baselines.
Appendix C Implementation Details
To facilitate reproducibility, we provide the complete source code for constructing the EEG–report benchmark from the Harvard Electroencephalography Database and the EHR database hosted on the Brain Data Science Platform, together with our model implementation, training scripts, data loaders, and pretrained weights at https://anonymous.4open.science/r/CELM-3AF4. Additionally, we provide the detailed implementation details of our approach in this section and the prompts used for the baselines in Appendix B. The dataset utilized for this study is publicly available and can be accessed at https://bdsp.io/content/harvard-eeg-db/4.1/ after obtaining credentialed access.
C.1 Model Details and Hyperparameters
We use CBraMod (Wang et al., 2024b) as the EEG encoder and Qwen-4B-Instruct-2507 as the local LLM backbone in our framework. CBraMod is selected due to its SOTA performance among EEG foundation models and its pretraining on large, diverse EEG datasets. In addition, our encoder ablation study comparing CBraMod with LaBraM shows that CBraMod consistently achieves better performance. We chose Qwen-4B-Instruct-2507 as the LLM backbone because it outperforms other local LLMs in unimodal baselines and reliably follows instruction prompts to generate outputs with the required structure. In Table 5, we provide the hyperparameters of the alignment modules used in our study.
Table 5: Hyperparameters of the alignment (projector) modules.
Module Hyperparameter Values
Linear Projector Projector Input Dimension (CBraMod Embedding Dimension) 200
Projector Output Dimension (Qwen-3 4B Embedding Dimension) 2560
Bias True
# Trainable parameters 522,240
Perceiver Projector # Query tokens $256$
Embedding Dimension 200
# Attention heads 88
# Perceiver layers (cross attention followed by feed-forward layer) 22
Dropout 0.1
Feed-forward multiplier ($\times$) 2
Final linear projection Linear(200, 2560)
# Trainable parameters 1,219,040
Sequence Context Compression (SCC) Transformer Encoder Type Linear attention transformer
Transformer # heads 88
Transformer # layers (depth) 11
# Query tokens $256$
# Perceiver layers 11
# Trainable parameters 1,378,440
Sequence Context Alignment (SCA) Transformer Encoder Type Linear attention transformer
# Heads 88
# Transformer Layers 22
Final linear projection Linear(200, 2560)
# Trainable parameters 1,486,240
Table 6: Training details
Hyperparameter Values
Batch size 4
Gradient accumulation steps 4
Optimizer AdamW
Learning rate 1e-4
Weight decay 0.01
$\beta_{1}$ 0.9
$\beta_{2}$ 0.99
LR scheduler Linear scheduler with warmup
# of training epochs 10
Warm-up ratio 0.1
Mixed precision bf16
C.2 Training Details
In our approach, the EEG encoder and LLM backbone are frozen, and only the alignment module is trained via supervised learning using a next-token prediction objective. Training details are summarized in Table 6.
C.3 Prompts
Figure 8 illustrates the prompt used in our approach
along with the EEG tokens. We also provide
examples of the sections in the prompt.
[…]
Your task is to generate the specified sections (**SECTIONS TO BE GENERATED**) of a formal clinical EEG report using the above provided data of EEG recording sessions and the following information:
[…]
Return ONLY the following JSON structure, with no preamble, explanation, or markdown:
[…]
[SECTION_NAMES]
[…]
e.g. [’EEG DESCRIPTION/DETAILS’] PATIENT HISTORY AND EEG DESCRIPTION
[PATIENT_HISTORY_AND_EEG_DESCRIPTION]
[…]
e.g. age: 77.0, gender: Female, indication: ***** y.o. female with history of afib, l mca stroke, and headaches, presenting with episodes of altered consciousness concerning for seizure vs. syncope. none pertinent medications: keppra, neurontin, seroquel. none Now generate the EEG report.
Figure 8: Prompt for CELM.
Appendix D Additional Experiment Results
D.1 Extended Report Generation Performance on S0001
Table 7 presents a comprehensive performance comparison of EEG report generation on the S0001 site for samples with clinical context. Both proposed variants consistently outperform all baselines across all evaluation metrics. Among them, CELM achieves better performance compared to the memory-efficient CELM -SCC variant across all metrics.
Table 7: Report generation performance comparison on samples with clinical context on S0001. We provide results for a complete set of evaluation metrics under two input settings: (i) Unimodal + Text Only Input, where language models generate reports solely from clinical context text, and (ii) Unimodal + Text + EEG Features Input, where EEG-derived features are additionally provided. The table compares multiple strong baselines, including general-purpose and medical LLMs, against our CELM . The best results are highlighted in orange, and the best baseline per category is highlighted in blue.
Method BLEU-1 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM BERTScore METEOR Unimodal + Text Only Input Gemma-3-1b-it 0.1897 $\pm$ 0.1340 0.0351 $\pm$ 0.0694 0.2112 $\pm$ 0.1408 0.0431 $\pm$ 0.0985 0.1356 $\pm$ 0.1129 0.1381 $\pm$ 0.1137 0.6461 $\pm$ 0.3014 0.1567 $\pm$ 0.1238 Gemma-3n-e2b-it 0.1967 $\pm$ 0.1592 0.0641 $\pm$ 0.1402 0.2410 $\pm$ 0.1722 0.0702 $\pm$ 0.1939 0.1657 $\pm$ 0.1826 0.1687 $\pm$ 0.1824 0.7848 $\pm$ 0.0765 0.1875 $\pm$ 0.1725 Gemma-3-4b-it 0.2372 $\pm$ 0.1508 0.0641 $\pm$ 0.1349 0.2857 $\pm$ 0.1590 0.0821 $\pm$ 0.1893 0.1893 $\pm$ 0.1749 0.1929 $\pm$ 0.1747 0.7941 $\pm$ 0.0922 0.2147 $\pm$ 0.1668 Gemma-3n-e4b-it 0.1793 $\pm$ 0.1685 0.0642 $\pm$ 0.1486 0.2416 $\pm$ 0.1778 0.0763 $\pm$ 0.2003 0.1659 $\pm$ 0.1892 0.1687 $\pm$ 0.1890 0.7847 $\pm$ 0.0747 0.1816 $\pm$ 0.1786 Medgemma-4b-it 0.1502 $\pm$ 0.1725 0.0485 $\pm$ 0.1383 0.2196 $\pm$ 0.1924 0.0654 $\pm$ 0.1954 0.1569 $\pm$ 0.1908 0.1593 $\pm$ 0.1910 0.6683 $\pm$ 0.2994 0.1529 $\pm$ 0.1835 Llama-3.2-1b-instruct 0.0337 $\pm$ 0.0667 0.0059 $\pm$ 0.0166 0.0919 $\pm$ 0.0744 0.0053 $\pm$ 0.0217 0.0615 $\pm$ 0.0500 0.0629 $\pm$ 0.0515 0.5580 $\pm$ 0.3284 0.0529 $\pm$ 0.0519 Llama-3.2-3b-instruct 0.1951 $\pm$ 0.1630 0.0625 $\pm$ 0.1487 0.2224 $\pm$ 0.1755 0.0669 $\pm$ 0.1860 0.1511 $\pm$ 0.1760 0.1550 $\pm$ 0.1759 0.7116 $\pm$ 0.2311 0.1843 $\pm$ 0.1864 Llama-3.1-8b-instruct 0.2128 $\pm$ 0.1702 0.0598 $\pm$ 0.1444 0.2511 $\pm$ 0.1872 0.0753 $\pm$ 0.1967 0.1689 $\pm$ 0.1889 0.1722 $\pm$ 0.1892 0.6994 $\pm$ 0.2605 0.1839 $\pm$ 0.1788 Meta-llama-3-8b-instruct 0.2249 $\pm$ 0.1602 0.0620 $\pm$ 0.1418 0.2673 $\pm$ 0.1731 0.0765 $\pm$ 0.1895 0.1758 $\pm$ 0.1785 0.1795 $\pm$ 0.1787 0.7392 $\pm$ 0.2055 0.1946 $\pm$ 0.1809 Qwen3-1.7b 0.1335 $\pm$ 0.1481 0.0213 $\pm$ 0.0450 0.1378 $\pm$ 0.1557 0.0285 $\pm$ 0.0704 0.0835 $\pm$ 0.1054 0.0855 $\pm$ 0.1075 0.3856 $\pm$ 0.3941 0.1102 $\pm$ 0.1316 Qwen3-4b-instruct-2507 0.2795 $\pm$ 0.1243 0.0593 $\pm$ 0.1100 0.3038 $\pm$ 0.1460 0.0763 $\pm$ 0.1672 0.1872 $\pm$ 0.1623 0.1916 $\pm$ 0.1622 0.7975 $\pm$ 0.0852 0.2418 $\pm$ 0.1440 Unimodal + Text + EEG Features Input Gemma-3-1b-it 0.0068 $\pm$ 0.0442 0.0019 $\pm$ 0.0273 0.0086 $\pm$ 0.0524 0.0022 $\pm$ 0.0367 0.0063 $\pm$ 0.0449 0.0064 $\pm$ 0.0451 0.0311 $\pm$ 0.1512 0.0064 $\pm$ 0.0462 Gemma-3n-e2b-it 0.0083 $\pm$ 0.0504 0.0029 $\pm$ 0.0330 0.0116 $\pm$ 0.0605 0.0032 $\pm$ 0.0403 0.0080 $\pm$ 0.0500 0.0083 $\pm$ 0.0507 0.0415 $\pm$ 0.1756 0.0088 $\pm$ 0.0543 Gemma-3-4b-it 0.0085 $\pm$ 0.0532 0.0025 $\pm$ 0.0299 0.0093 $\pm$ 0.0594 0.0032 $\pm$ 0.0402 0.0062 $\pm$ 0.0466 0.0065 $\pm$ 0.0474 0.0247 $\pm$ 0.1384 0.0078 $\pm$ 0.0558 Gemma-3n-e4b-it 0.0101 $\pm$ 0.0348 0.0020 $\pm$ 0.0067 0.0397 $\pm$ 0.0606 0.0014 $\pm$ 0.0072 0.0259 $\pm$ 0.0395 0.0262 $\pm$ 0.0401 0.2498 $\pm$ 0.3570 0.0231 $\pm$ 0.0392 Medgemma-4b-it 0.1618 $\pm$ 0.1730 0.0488 $\pm$ 0.1400 0.2100 $\pm$ 0.1919 0.0611 $\pm$ 0.1870 0.1455 $\pm$ 0.1858 0.1482 $\pm$ 0.1862 0.6172 $\pm$ 0.3319 0.1529 $\pm$ 0.1832 Llama-3.2-1b-instruct 0.0310 $\pm$ 0.0615 0.0053 $\pm$ 0.0119 0.0820 $\pm$ 0.0727 0.0037 $\pm$ 0.0150 0.0547 $\pm$ 0.0485 0.0556 $\pm$ 0.0493 0.4978 $\pm$ 0.3562 0.0473 $\pm$ 0.0479 Llama-3.2-3b-instruct 0.1692 $\pm$ 0.1414 0.0528 $\pm$ 0.1147 0.1844 $\pm$ 0.1611 0.0543 $\pm$ 0.1575 0.1271 $\pm$ 0.1529 0.1342 $\pm$ 0.1536 0.6441 $\pm$ 0.2898 0.1751 $\pm$ 0.1871 Llama-3.1-8b-instruct 0.2040 $\pm$ 0.1450 0.0508 $\pm$ 0.1177 0.2151 $\pm$ 0.1693 0.0617 $\pm$ 0.1695 0.1469 $\pm$ 0.1690 0.1499 $\pm$ 0.1693 0.6930 $\pm$ 0.2289 0.1844 $\pm$ 0.1647 Qwen3-1.7b 0.2047 $\pm$ 0.1351 0.0331 $\pm$ 0.0611 0.2051 $\pm$ 0.1453 0.0422 $\pm$ 0.0971 0.1286 $\pm$ 0.1150 0.1317 $\pm$ 0.1163 0.6063 $\pm$ 0.3215 0.1683 $\pm$ 0.1304 Qwen3-4b-instruct-2507 0.1809 $\pm$ 0.0986 0.0288 $\pm$ 0.0555 0.2058 $\pm$ 0.1192 0.0402 $\pm$ 0.0969 0.1275 $\pm$ 0.1082 0.1303 $\pm$ 0.1087 0.7249 $\pm$ 0.1840 0.2067 $\pm$ 0.1079 CELM-SCC 0.3383 $\pm$ 0.1936 0.1345 $\pm$ 0.1860 0.3843 $\pm$ 0.1876 0.1651 $\pm$ 0.1960 0.2699 $\pm$ 0.1867 0.2732 $\pm$ 0.1865 0.7949 $\pm$ 0.1783 0.2889 $\pm$ 0.1866 CELM 0.4823 $\pm$ 0.1920 0.2831 $\pm$ 0.1952 0.5565 $\pm$ 0.1683 0.3458 $\pm$ 0.2036 0.4687 $\pm$ 0.1857 0.4702 $\pm$ 0.1852 0.8697 $\pm$ 0.0901 0.4734 $\pm$ 0.1941
D.2 Extended Report Generation Performance on S0002
Tables 8 and 9 present a comprehensive performance comparison of EEG report generation on the S0002 site under settings with and without patient clinical context. Both proposed variants consistently outperform all baselines across all evaluation metrics. Among them, CELM achieves better performance compared to the memory-efficient CELM -SCC variant across all metrics, except for BERTScore in the zero-context setting.
Table 8: Report generation performance comparison on samples with clinical context on S0002. We provide results for a complete set of evaluation metrics under two input settings: (i) Unimodal + Text Only Input, where language models generate reports solely from clinical context text, and (ii) Unimodal + Text + EEG Features Input, where EEG-derived features are additionally provided. The table compares multiple strong baselines, including general-purpose and medical LLMs, against our CELM . The best results are highlighted in orange, and the best baseline per category is highlighted in blue.
Method BLEU-1 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM BERTScore METEOR Unimodal + Text Only Input Gemma-3-1b-it 0.1509 $\pm$ 0.0898 0.0158 $\pm$ 0.0121 0.2330 $\pm$ 0.0876 0.0302 $\pm$ 0.0210 0.1233 $\pm$ 0.0467 0.1259 $\pm$ 0.0495 0.7058 $\pm$ 0.2423 0.1319 $\pm$ 0.0565 Gemma-3n-e2b-it 0.0925 $\pm$ 0.0769 0.0115 $\pm$ 0.0230 0.2318 $\pm$ 0.0600 0.0274 $\pm$ 0.0346 0.1224 $\pm$ 0.0377 0.1254 $\pm$ 0.0408 0.7683 $\pm$ 0.0560 0.1181 $\pm$ 0.0472 Gemma-3-4b-it 0.1538 $\pm$ 0.0990 0.0163 $\pm$ 0.0180 0.2775 $\pm$ 0.0558 0.0354 $\pm$ 0.0335 0.1407 $\pm$ 0.0336 0.1440 $\pm$ 0.0369 0.7851 $\pm$ 0.0730 0.1506 $\pm$ 0.0514 Gemma-3n-e4b-it 0.0539 $\pm$ 0.0664 0.0083 $\pm$ 0.0243 0.1998 $\pm$ 0.0649 0.0281 $\pm$ 0.0358 0.1152 $\pm$ 0.0392 0.1175 $\pm$ 0.0407 0.7635 $\pm$ 0.0525 0.0993 $\pm$ 0.0490 Medgemma-4b-it 0.0865 $\pm$ 0.0855 0.0109 $\pm$ 0.0194 0.2280 $\pm$ 0.0734 0.0341 $\pm$ 0.0341 0.1315 $\pm$ 0.0433 0.1345 $\pm$ 0.0458 0.7576 $\pm$ 0.1375 0.1198 $\pm$ 0.0542 Llama-3.2-1b-instruct 0.0340 $\pm$ 0.0700 0.0036 $\pm$ 0.0073 0.0954 $\pm$ 0.0869 0.0082 $\pm$ 0.0135 0.0581 $\pm$ 0.0483 0.0596 $\pm$ 0.0509 0.5139 $\pm$ 0.3501 0.0494 $\pm$ 0.0504 Llama-3.2-3b-instruct 0.1651 $\pm$ 0.1110 0.0186 $\pm$ 0.0256 0.2346 $\pm$ 0.0877 0.0321 $\pm$ 0.0351 0.1221 $\pm$ 0.0478 0.1271 $\pm$ 0.0517 0.7261 $\pm$ 0.1919 0.1430 $\pm$ 0.0689 Llama-3.1-8b-instruct 0.1629 $\pm$ 0.0970 0.0190 $\pm$ 0.0237 0.2655 $\pm$ 0.0766 0.0424 $\pm$ 0.0356 0.1368 $\pm$ 0.0444 0.1404 $\pm$ 0.0476 0.7502 $\pm$ 0.1531 0.1480 $\pm$ 0.0581 Meta-llama-3-8b-instruct 0.1441 $\pm$ 0.0847 0.0176 $\pm$ 0.0232 0.2588 $\pm$ 0.0586 0.0403 $\pm$ 0.0330 0.1354 $\pm$ 0.0374 0.1390 $\pm$ 0.0411 0.7702 $\pm$ 0.0817 0.1416 $\pm$ 0.0481 Qwen3-1.7b 0.0886 $\pm$ 0.1095 0.0097 $\pm$ 0.0124 0.1189 $\pm$ 0.1296 0.0142 $\pm$ 0.0191 0.0633 $\pm$ 0.0686 0.0652 $\pm$ 0.0716 0.3650 $\pm$ 0.3893 0.0735 $\pm$ 0.0821 Qwen3-4b-instruct-2507 0.1912 $\pm$ 0.0807 0.0218 $\pm$ 0.0197 0.2690 $\pm$ 0.0448 0.0398 $\pm$ 0.0298 0.1317 $\pm$ 0.0295 0.1350 $\pm$ 0.0324 0.7867 $\pm$ 0.0178 0.1622 $\pm$ 0.0434 Unimodal + Text + EEG Features Input Gemma-3-1b-it 0.0957 $\pm$ 0.0925 0.0109 $\pm$ 0.0203 0.1755 $\pm$ 0.1157 0.0239 $\pm$ 0.0316 0.0980 $\pm$ 0.0653 0.1007 $\pm$ 0.0684 0.5785 $\pm$ 0.3384 0.0954 $\pm$ 0.0692 Gemma-3n-e2b-it 0.0750 $\pm$ 0.0771 0.0093 $\pm$ 0.0226 0.2190 $\pm$ 0.0840 0.0350 $\pm$ 0.0350 0.1256 $\pm$ 0.0516 0.1281 $\pm$ 0.0539 0.7175 $\pm$ 0.2153 0.1082 $\pm$ 0.0546 Gemma-3-4b-it 0.2089 $\pm$ 0.0895 0.0222 $\pm$ 0.0186 0.2886 $\pm$ 0.0458 0.0446 $\pm$ 0.0331 0.1451 $\pm$ 0.0315 0.1491 $\pm$ 0.0356 0.7883 $\pm$ 0.0418 0.1738 $\pm$ 0.0476 Gemma-3n-e4b-it 0.0169 $\pm$ 0.0486 0.0018 $\pm$ 0.0052 0.0357 $\pm$ 0.0903 0.0049 $\pm$ 0.0138 0.0194 $\pm$ 0.0490 0.0207 $\pm$ 0.0526 0.1073 $\pm$ 0.2678 0.0180 $\pm$ 0.0462 Medgemma-4b-it 0.0888 $\pm$ 0.0910 0.0110 $\pm$ 0.0195 0.1829 $\pm$ 0.1128 0.0248 $\pm$ 0.0349 0.1008 $\pm$ 0.0641 0.1029 $\pm$ 0.0660 0.5979 $\pm$ 0.3260 0.1000 $\pm$ 0.0702 Llama-3.2-1b-instruct 0.0312 $\pm$ 0.0668 0.0035 $\pm$ 0.0073 0.1016 $\pm$ 0.0773 0.0090 $\pm$ 0.0142 0.0620 $\pm$ 0.0420 0.0634 $\pm$ 0.0438 0.5903 $\pm$ 0.3062 0.0517 $\pm$ 0.0448 Llama-3.2-3b-instruct 0.1779 $\pm$ 0.1101 0.0207 $\pm$ 0.0224 0.2309 $\pm$ 0.0903 0.0350 $\pm$ 0.0349 0.1217 $\pm$ 0.0484 0.1364 $\pm$ 0.0575 0.7377 $\pm$ 0.1557 0.1540 $\pm$ 0.0743 Llama-3.1-8b-instruct 0.1960 $\pm$ 0.0959 0.0227 $\pm$ 0.0251 0.2256 $\pm$ 0.0904 0.0365 $\pm$ 0.0364 0.1198 $\pm$ 0.0542 0.1236 $\pm$ 0.0574 0.6895 $\pm$ 0.2171 0.1509 $\pm$ 0.0682 Qwen3-1.7b 0.1890 $\pm$ 0.1048 0.0185 $\pm$ 0.0118 0.2049 $\pm$ 0.1021 0.0236 $\pm$ 0.0179 0.1077 $\pm$ 0.0530 0.1110 $\pm$ 0.0568 0.6345 $\pm$ 0.2944 0.1413 $\pm$ 0.0724 Qwen3-4b-instruct-2507 0.2483 $\pm$ 0.0425 0.0217 $\pm$ 0.0098 0.2638 $\pm$ 0.0341 0.0334 $\pm$ 0.0179 0.1283 $\pm$ 0.0180 0.1322 $\pm$ 0.0250 0.7704 $\pm$ 0.0125 0.2017 $\pm$ 0.0285 CELM-SCC 0.3767 $\pm$ 0.1557 0.1671 $\pm$ 0.1054 0.4487 $\pm$ 0.1283 0.2149 $\pm$ 0.1109 0.3232 $\pm$ 0.1263 0.3268 $\pm$ 0.1258 0.8229 $\pm$ 0.1341 $32±\pm 0$.1261 CELM 0.5695 $\pm$ 0.1702 0.4145 $\pm$ 0.1639 0.6408 $\pm$ 0.1494 0.4805 $\pm$ 0.1603 0.5757 $\pm$ 0.1624 0.5772 $\pm$ 0.1613 0.8755 $\pm$ 0.1521 0.5597 $\pm$ 0.1728
Table 9: Zero-context report generation performance on S0002. The best results are highlighted in orange, and the best baseline is highlighted in blue.
Method BLEU-1 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM BERTScore METEOR Unimodal + Text + EEG Features Input Gemma-3-1b-it 0.1019 $\pm$ 0.0782 0.0105 $\pm$ 0.0082 0.1935 $\pm$ 0.0953 0.0185 $\pm$ 0.0148 0.1046 $\pm$ 0.0509 0.1046 $\pm$ 0.0509 0.6441 $\pm$ 0.2920 0.1052 $\pm$ 0.0556 Gemma-3n-e2b-it 0.0717 $\pm$ 0.0710 0.0089 $\pm$ 0.0090 0.2203 $\pm$ 0.0581 0.0331 $\pm$ 0.0175 0.1341 $\pm$ 0.0331 0.1346 $\pm$ 0.0332 0.7716 $\pm$ 0.0826 0.1116 $\pm$ 0.0411 Gemma-3-4b-it 0.1699 $\pm$ 0.0885 0.0189 $\pm$ 0.0122 0.2557 $\pm$ 0.0484 0.0382 $\pm$ 0.0203 0.1341 $\pm$ 0.0271 0.1341 $\pm$ 0.0271 0.7875 $\pm$ 0.0203 0.1514 $\pm$ 0.0412 Gemma-3n-e4b-it 0.0058 $\pm$ 0.0304 0.0007 $\pm$ 0.0037 0.0127 $\pm$ 0.0509 0.0021 $\pm$ 0.0094 0.0072 $\pm$ 0.0288 0.0072 $\pm$ 0.0288 0.0471 $\pm$ 0.1844 0.0067 $\pm$ 0.0281 Medgemma-4b-it 0.0559 $\pm$ 0.0660 0.0069 $\pm$ 0.0080 0.1591 $\pm$ 0.0959 0.0208 $\pm$ 0.0191 0.0917 $\pm$ 0.0535 0.0917 $\pm$ 0.0535 0.6073 $\pm$ 0.3121 0.0820 $\pm$ 0.0520 Llama-3.2-1b-instruct 0.0265 $\pm$ 0.0627 0.0033 $\pm$ 0.0074 0.1041 $\pm$ 0.0723 0.0109 $\pm$ 0.0116 0.0632 $\pm$ 0.0361 0.0632 $\pm$ 0.0361 0.6324 $\pm$ 0.2729 0.0530 $\pm$ 0.0396 Llama-3.2-3b-instruct 0.1261 $\pm$ 0.1028 0.0125 $\pm$ 0.0094 0.1995 $\pm$ 0.0806 0.0304 $\pm$ 0.0188 0.1114 $\pm$ 0.0425 0.1141 $\pm$ 0.0453 0.7328 $\pm$ 0.1553 0.1214 $\pm$ 0.0617 Llama-3.1-8b-instruct 0.1556 $\pm$ 0.0817 0.0160 $\pm$ 0.0088 0.2090 $\pm$ 0.0737 0.0305 $\pm$ 0.0170 0.1142 $\pm$ 0.0416 0.1149 $\pm$ 0.0417 0.7010 $\pm$ 0.1985 0.1298 $\pm$ 0.0506 Qwen3-1.7b 0.1772 $\pm$ 0.0999 0.0164 $\pm$ 0.0104 0.1819 $\pm$ 0.0938 0.0187 $\pm$ 0.0149 0.0982 $\pm$ 0.0498 0.0991 $\pm$ 0.0502 0.6265 $\pm$ 0.3013 0.1304 $\pm$ 0.0688 Qwen3-4b-instruct-2507 0.2260 $\pm$ 0.0592 0.0183 $\pm$ 0.0071 0.2315 $\pm$ 0.0501 0.0241 $\pm$ 0.0139 0.1162 $\pm$ 0.0254 0.1162 $\pm$ 0.0254 0.7418 $\pm$ 0.1337 0.1790 $\pm$ 0.0420 CELM-SCC 0.2991 $\pm$ 0.1276 0.0912 $\pm$ 0.0590 0.3793 $\pm$ 0.0892 0.1378 $\pm$ 0.0666 0.2441 $\pm$ 0.0771 0.2441 $\pm$ 0.0771 0.8192 $\pm$ 0.0299 0.2574 $\pm$ 0.0891 CELM 0.4652 $\pm$ 0.1884 0.2666 $\pm$ 0.1354 0.5248 $\pm$ 0.1869 0.3271 $\pm$ 0.1494 0.4339 $\pm$ 0.1767 0.4339 $\pm$ 0.1767 0.7990 $\pm$ 0.2432 0.4390 $\pm$ 0.1869
D.3 Additional Results on Alignment Module Ablation
Comprehensive results of the alignment-module ablation study on the S0002 dataset, covering settings with clinical context, without clinical context, and all cases, are summarized in Table 10. Across most evaluation settings, the SCT projector achieves significantly stronger performance than other alignment variants, highlighting the importance of modeling dependencies among EEG epoch tokens before projecting them into the LLM embedding space.
Table 10: Alignment module ablation on S0002 dataset. The best results are highlighted in orange
Projector BLEU-1 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM BERTScore METEOR EEG Only Linear Projector 0.2644 $\pm$ 0.0471 0.0247 $\pm$ 0.0063 0.2928 $\pm$ 0.0314 0.0447 $\pm$ 0.0141 0.1392 $\pm$ 0.0161 0.1392 $\pm$ 0.0161 0.7828 $\pm$ 0.0152 0.2072 $\pm$ 0.0382 Perceiver Projector 0.2636 $\pm$ 0.0696 0.0249 $\pm$ 0.0098 0.2897 $\pm$ 0.0344 0.0404 $\pm$ 0.0146 0.1377 $\pm$ 0.0165 0.1377 $\pm$ 0.0165 0.7888 $\pm$ 0.0162 0.1968 $\pm$ 0.0385 SCT Perceiver Projector 0.2991 $\pm$ 0.1276 0.0912 $\pm$ 0.0590 0.3793 $\pm$ 0.0892 0.1378 $\pm$ 0.0666 0.2441 $\pm$ 0.0771 0.2441 $\pm$ 0.0771 0.8192 $\pm$ 0.0299 0.2574 $\pm$ 0.0891 SCT Projector 0.4652 $\pm$ 0.1884 0.2666 $\pm$ 0.1354 0.5248 $\pm$ 0.1869 0.3271 $\pm$ 0.1494 0.4339 $\pm$ 0.1767 0.4339 $\pm$ 0.1767 0.7990 $\pm$ 0.2432 0.4390 $\pm$ 0.1869 With Patient History Linear Projector 0.2775 $\pm$ 0.0485 0.0283 $\pm$ 0.0121 0.2998 $\pm$ 0.0425 0.0507 $\pm$ 0.0240 0.1429 $\pm$ 0.0239 0.1473 $\pm$ 0.0306 0.7852 $\pm$ 0.0153 0.2143 $\pm$ 0.0349 Perceiver Projector 0.2672 $\pm$ 0.0556 0.0273 $\pm$ 0.0192 0.2879 $\pm$ 0.0378 0.0461 $\pm$ 0.0300 0.1384 $\pm$ 0.0281 0.1422 $\pm$ 0.0321 0.7878 $\pm$ 0.0159 0.1955 $\pm$ 0.0393 SCT Perceiver Projector 0.3767 $\pm$ 0.1557 0.1671 $\pm$ 0.1054 0.4487 $\pm$ 0.1283 0.2149 $\pm$ 0.1109 0.3232 $\pm$ 0.1263 0.3268 $\pm$ 0.1258 0.8229 $\pm$ 0.1341 0.3232 $\pm$ 0.1261 SCT Projector 0.5695 $\pm$ 0.1702 0.4145 $\pm$ 0.1639 0.6408 $\pm$ 0.1494 0.4805 $\pm$ 0.1603 0.5757 $\pm$ 0.1624 0.5772 $\pm$ 0.1613 0.8755 $\pm$ 0.1521 0.5597 $\pm$ 0.1728 All Linear Projector 0.2762 $\pm$ 0.0485 0.0279 $\pm$ 0.0117 0.2991 $\pm$ 0.0416 0.0501 $\pm$ 0.0233 0.1426 $\pm$ 0.0232 0.1465 $\pm$ 0.0296 0.7849 $\pm$ 0.0153 0.2135 $\pm$ 0.0353 Perceiver Projector 0.2669 $\pm$ 0.0571 0.0271 $\pm$ 0.0184 0.2881 $\pm$ 0.0375 0.0455 $\pm$ 0.0288 0.1384 $\pm$ 0.0271 0.1417 $\pm$ 0.0309 0.7879 $\pm$ 0.0160 0.1956 $\pm$ 0.0392 SCT Perceiver Projector 0.3689 $\pm$ 0.1548 0.1594 $\pm$ 0.1042 0.4417 $\pm$ 0.1267 0.2071 $\pm$ 0.1098 0.3152 $\pm$ 0.1246 0.3185 $\pm$ 0.1243 0.8226 $\pm$ 0.1275 0.3166 $\pm$ 0.1245 SCT Projector 0.5590 $\pm$ 0.1749 0.3996 $\pm$ 0.1673 0.6292 $\pm$ 0.1574 0.4650 $\pm$ 0.1657 0.5614 $\pm$ 0.1692 0.5628 $\pm$ 0.1685 0.8678 $\pm$ 0.1650 0.5475 $\pm$ 0.1779
D.4 EEG Encoder Ablation
We conducted an ablation study to analyze report generation performance across different EEG encoders. Specifically, we compared CBraMod (Wang et al., 2024b) and LaBraM (Jiang et al., 2024), both EEG foundation models pretrained on large-scale EEG datasets. The results show that CBraMod consistently outperforms LaBraM, highlighting the critical role of high-quality EEG representations in effective clinical report generation.
Table 11: EEG encoder ablation study
EEG Encoder BLEU-1 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM BERTScore METEOR Labram 0.4654 $\pm$ 0.2272 0.3188 $\pm$ 0.1904 0.5355 $\pm$ 0.2312 0.3809 $\pm$ 0.2027 0.4673 $\pm$ 0.2233 0.4689 $\pm$ 0.2231 0.7781 $\pm$ 0.2931 0.4600 $\pm$ 0.2259 Cbramod 0.5590 $\pm$ 0.1749 0.3996 $\pm$ 0.1673 0.6292 $\pm$ 0.1574 0.4650 $\pm$ 0.1657 0.5614 $\pm$ 0.1692 0.5628 $\pm$ 0.1685 0.8678 $\pm$ 0.1650 0.5475 $\pm$ 0.1779
D.5 Score Distribution
Figure 9: Distribution of various metrics for overall report generation in the S0002 dataset.
Figure 10: Distribution of various metrics for overall report generation in the S0001 dataset.
D.6 More Results on Section-wise Analysis
Table 12: Performance by report section on S0001 dataset
Report Section Method BLEU-1 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM METEOR EEG Description/Details gemma-3n-E4B-it 0.0095 $\pm$ 0.0344 0.0017 $\pm$ 0.0057 0.0396 $\pm$ 0.0604 0.0012 $\pm$ 0.0057 0.0256 $\pm$ 0.0389 0.0259 $\pm$ 0.0395 0.0228 $\pm$ 0.0385 medgemma-4b-it 0.1624 $\pm$ 0.1729 0.0489 $\pm$ 0.1401 0.2108 $\pm$ 0.1917 0.0612 $\pm$ 0.1872 0.1460 $\pm$ 0.1858 0.1487 $\pm$ 0.1861 0.1535 $\pm$ 0.1831 Llama-3.1-8B-Instruct 0.2045 $\pm$ 0.1452 0.0501 $\pm$ 0.1179 0.2143 $\pm$ 0.1691 0.0603 $\pm$ 0.1694 0.1454 $\pm$ 0.1687 0.1485 $\pm$ 0.1690 0.1838 $\pm$ 0.1645 Qwen3-4B-Instruct-2507 0.1818 $\pm$ 0.0988 0.0285 $\pm$ 0.0555 0.2059 $\pm$ 0.1190 0.0393 $\pm$ 0.0969 0.1271 $\pm$ 0.1082 0.1300 $\pm$ 0.1087 0.2069 $\pm$ 0.1080 CELM-SCC 0.3381 $\pm$ 0.1900 0.1291 $\pm$ 0.1832 0.3826 $\pm$ 0.1854 0.1601 $\pm$ 0.1935 0.2650 $\pm$ 0.1841 0.2687 $\pm$ 0.1840 0.2875 $\pm$ 0.1833 CELM 0.4907 $\pm$ 0.1834 0.2841 $\pm$ 0.1911 0.5631 $\pm$ 0.1607 0.3483 $\pm$ 0.1987 0.4727 $\pm$ 0.1813 0.4745 $\pm$ 0.1809 0.4798 $\pm$ 0.1873 Epileptiform Abnormalities gemma-3n-E4B-it 0.0212 $\pm$ 0.0508 0.0081 $\pm$ 0.0201 0.0256 $\pm$ 0.0642 0.0048 $\pm$ 0.0181 0.0212 $\pm$ 0.0529 0.0217 $\pm$ 0.0539 0.0193 $\pm$ 0.0563 medgemma-4b-it 0.0015 $\pm$ 0.0117 0.0008 $\pm$ 0.0064 0.0043 $\pm$ 0.0330 0.0034 $\pm$ 0.0265 0.0043 $\pm$ 0.0330 0.0043 $\pm$ 0.0330 0.0030 $\pm$ 0.0234 Llama-3.1-8B-Instruct 0.0786 $\pm$ 0.1185 0.0343 $\pm$ 0.0596 0.0954 $\pm$ 0.1437 0.0422 $\pm$ 0.0860 0.0858 $\pm$ 0.1337 0.0863 $\pm$ 0.1347 0.0992 $\pm$ 0.1523 Qwen3-4B-Instruct-2507 0.0574 $\pm$ 0.0919 0.0150 $\pm$ 0.0240 0.0772 $\pm$ 0.1201 0.0273 $\pm$ 0.0473 0.0568 $\pm$ 0.0808 0.0578 $\pm$ 0.0830 0.0759 $\pm$ 0.0942 CELM-SCC 0.2870 $\pm$ 0.3051 0.1862 $\pm$ 0.2173 0.3392 $\pm$ 0.2614 0.2104 $\pm$ 0.2319 0.3233 $\pm$ 0.2597 0.3221 $\pm$ 0.2598 0.2717 $\pm$ 0.2560 CELM 0.3172 $\pm$ 0.3472 0.2061 $\pm$ 0.2641 0.3881 $\pm$ 0.2967 0.2401 $\pm$ 0.2796 0.3600 $\pm$ 0.2902 0.3607 $\pm$ 0.2897 0.3222 $\pm$ 0.3107 Background Activity gemma-3n-E4B-it 0.0440 $\pm$ 0.0909 0.0058 $\pm$ 0.0113 0.0954 $\pm$ 0.1133 0.0071 $\pm$ 0.0175 0.0584 $\pm$ 0.0679 0.0584 $\pm$ 0.0679 0.0534 $\pm$ 0.0685 medgemma-4b-it 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 Llama-3.1-8B-Instruct 0.0318 $\pm$ 0.0706 0.0047 $\pm$ 0.0102 0.1093 $\pm$ 0.1004 0.0102 $\pm$ 0.0181 0.0759 $\pm$ 0.0690 0.0759 $\pm$ 0.0690 0.0443 $\pm$ 0.0433 Qwen3-4B-Instruct-2507 0.2172 $\pm$ 0.1394 0.0302 $\pm$ 0.0202 0.1984 $\pm$ 0.1289 0.0396 $\pm$ 0.0306 0.1158 $\pm$ 0.0752 0.1158 $\pm$ 0.0752 0.1492 $\pm$ 0.0969 CELM-SCC 0.1494 $\pm$ 0.1142 0.0266 $\pm$ 0.0281 0.2442 $\pm$ 0.0629 0.0441 $\pm$ 0.0252 0.1670 $\pm$ 0.0521 0.1670 $\pm$ 0.0521 0.1355 $\pm$ 0.0524 CELM 0.1825 $\pm$ 0.1225 0.0332 $\pm$ 0.0244 0.2703 $\pm$ 0.0536 0.0701 $\pm$ 0.0476 0.1935 $\pm$ 0.0530 0.1935 $\pm$ 0.0530 0.1786 $\pm$ 0.0655 Interictal Epileptiform Abnormalities gemma-3n-E4B-it 0.0207 $\pm$ 0.0414 0.0096 $\pm$ 0.0192 0.0255 $\pm$ 0.0318 0.0075 $\pm$ 0.0150 0.0255 $\pm$ 0.0318 0.0255 $\pm$ 0.0318 0.0159 $\pm$ 0.0182 medgemma-4b-it 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 Llama-3.1-8B-Instruct 0.0891 $\pm$ 0.0899 0.0125 $\pm$ 0.0126 0.1200 $\pm$ 0.1008 0.0228 $\pm$ 0.0221 0.0672 $\pm$ 0.0550 0.0708 $\pm$ 0.0581 0.0863 $\pm$ 0.0781 Qwen3-4B-Instruct-2507 0.1594 $\pm$ 0.0755 0.0251 $\pm$ 0.0092 0.1856 $\pm$ 0.1088 0.0334 $\pm$ 0.0218 0.0985 $\pm$ 0.0502 0.1048 $\pm$ 0.0539 0.1253 $\pm$ 0.0472 CELM-SCC 0.0100 $\pm$ 0.0133 0.0027 $\pm$ 0.0039 0.0739 $\pm$ 0.0444 0.0228 $\pm$ 0.0140 0.0635 $\pm$ 0.0361 0.0666 $\pm$ 0.0338 0.0457 $\pm$ 0.0217 CELM 0.0765 $\pm$ 0.0815 0.0135 $\pm$ 0.0129 0.1166 $\pm$ 0.1240 0.0236 $\pm$ 0.0232 0.0787 $\pm$ 0.0779 0.0787 $\pm$ 0.0779 0.0650 $\pm$ 0.0559 Events/Seizures gemma-3n-E4B-it 0.0316 $\pm$ 0.0632 0.0086 $\pm$ 0.0172 0.0497 $\pm$ 0.0696 0.0012 $\pm$ 0.0016 0.0286 $\pm$ 0.0341 0.0291 $\pm$ 0.0341 0.0364 $\pm$ 0.0610 medgemma-4b-it 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 Llama-3.1-8B-Instruct 0.0448 $\pm$ 0.0551 0.0225 $\pm$ 0.0302 0.0544 $\pm$ 0.0495 0.0124 $\pm$ 0.0211 0.0528 $\pm$ 0.0497 0.0528 $\pm$ 0.0497 0.0212 $\pm$ 0.0201 Qwen3-4B-Instruct-2507 0.0511 $\pm$ 0.0465 0.0167 $\pm$ 0.0138 0.0955 $\pm$ 0.0374 0.0040 $\pm$ 0.0071 0.0749 $\pm$ 0.0315 0.0758 $\pm$ 0.0300 0.0677 $\pm$ 0.0648 CELM-SCC 0.0481 $\pm$ 0.0674 0.0257 $\pm$ 0.0353 0.1617 $\pm$ 0.1375 0.0004 $\pm$ 0.0009 0.1571 $\pm$ 0.1414 0.1578 $\pm$ 0.1407 0.0290 $\pm$ 0.0309 CELM 0.0110 $\pm$ 0.0220 0.0090 $\pm$ 0.0181 0.1606 $\pm$ 0.1248 0.0024 $\pm$ 0.0033 0.1512 $\pm$ 0.1322 0.1531 $\pm$ 0.1305 0.0288 $\pm$ 0.0119 Seizures gemma-3n-E4B-it 0.0168 $\pm$ 0.0352 0.0032 $\pm$ 0.0070 0.0405 $\pm$ 0.0588 0.0005 $\pm$ 0.0013 0.0256 $\pm$ 0.0343 0.0300 $\pm$ 0.0409 0.0187 $\pm$ 0.0273 medgemma-4b-it 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 Llama-3.1-8B-Instruct 0.0273 $\pm$ 0.0489 0.0169 $\pm$ 0.0328 0.0294 $\pm$ 0.0432 0.0010 $\pm$ 0.0026 0.0279 $\pm$ 0.0427 0.0294 $\pm$ 0.0432 0.0286 $\pm$ 0.0542 Qwen3-4B-Instruct-2507 0.0388 $\pm$ 0.0558 0.0121 $\pm$ 0.0180 0.0790 $\pm$ 0.1025 0.0085 $\pm$ 0.0149 0.0502 $\pm$ 0.0552 0.0544 $\pm$ 0.0606 0.0464 $\pm$ 0.0543 CELM-SCC 0.0631 $\pm$ 0.1319 0.0551 $\pm$ 0.1180 0.0773 $\pm$ 0.1057 0.0000 $\pm$ 0.0000 0.0725 $\pm$ 0.1075 0.0735 $\pm$ 0.1072 0.0336 $\pm$ 0.0502 CELM 0.0272 $\pm$ 0.0489 0.0253 $\pm$ 0.0471 0.1390 $\pm$ 0.2061 0.0005 $\pm$ 0.0012 0.1285 $\pm$ 0.2088 0.1329 $\pm$ 0.2081 0.0419 $\pm$ 0.0545 Impression/Interpretation gemma-3n-E4B-it 0.0172 $\pm$ 0.0725 0.0108 $\pm$ 0.0622 0.0237 $\pm$ 0.0949 0.0125 $\pm$ 0.0876 0.0207 $\pm$ 0.0923 0.0207 $\pm$ 0.0923 0.0193 $\pm$ 0.0990 medgemma-4b-it 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 Llama-3.1-8B-Instruct 0.0809 $\pm$ 0.1339 0.0294 $\pm$ 0.0774 0.1148 $\pm$ 0.1863 0.0489 $\pm$ 0.1334 0.0883 $\pm$ 0.1654 0.0895 $\pm$ 0.1663 0.0876 $\pm$ 0.1499 Qwen3-4B-Instruct-2507 0.0885 $\pm$ 0.1173 0.0201 $\pm$ 0.0328 0.1069 $\pm$ 0.1399 0.0308 $\pm$ 0.0570 0.0675 $\pm$ 0.0910 0.0693 $\pm$ 0.0945 0.0996 $\pm$ 0.1172 CELM-SCC 0.2194 $\pm$ 0.2899 0.1465 $\pm$ 0.2993 0.2687 $\pm$ 0.2946 0.1531 $\pm$ 0.3116 0.2190 $\pm$ 0.2929 0.2215 $\pm$ 0.2929 0.2206 $\pm$ 0.2996 CELM 0.2493 $\pm$ 0.3481 0.2103 $\pm$ 0.3515 0.3581 $\pm$ 0.3381 0.2473 $\pm$ 0.3793 0.3358 $\pm$ 0.3434 0.3372 $\pm$ 0.3428 0.2829 $\pm$ 0.3428
Table 12 reports the complete section-wise performance across all models and evaluation metrics. These results extend the main text by providing detailed comparisons for each report section and further show the consistency of CELM across diverse clinical report components.
D.7 Qualitative Analysis and Case Studies
This subsection presents qualitative analyses and representative case studies to complement the quantitative evaluation. In Figure 111213, and  14, we compare generated reports from the unimodal baseline, CELM-SCC, and CELM against ground-truth reports across multiple EEG report sections and datasets.
Figure 11: Impression/interpretation generation examples on S0001. Examples are ordered from (a) to (d) by decreasing ROUGE-1 score. We compare outputs from the unimodal baseline (text + EEG features), CELM-SCC, CELM, and the ground-truth reports.
Figure 12: Section-wise generation examples on S0001. (a) Epileptiform abnormalities, (b) Interictal epileptiform abnormalities, (c) Seizures, and (d) Background activity. We compare outputs from the unimodal baseline (text + EEG features), CELM-SCC, CELM, and the ground-truth reports.
Figure 13: EEG description/details generation examples on S0001. Examples are ordered from (a) to (d) by decreasing ROUGE-1 score. We compare outputs from the unimodal baseline (text + EEG features), CELM-SCC, CELM, and the ground-truth reports.
Figure 14: EEG description/details generation examples on S0002. Examples are ordered from (a) to (d) by decreasing ROUGE-1 score. We compare outputs from the unimodal baseline (text + EEG features), CELM-SCC, CELM, and the ground-truth reports.